Twice the first number is increased by the second number is 16. Find the two numbers if the sum of their squares is a minimum.
Pleassse help

Answer by Alan3354(69443)

Twice the first number is increased by the second number is 16. Find the two numbers if the sum of their squares is a minimum.
-------------
2x + y = 16
x^2 + y^2 = s
----------
y = 16 - 2x
x^2 + (16-2x)^2 = s
x^2 + 4x^2 - 64x + 256 = s
5x^2 - 64x + 256 = s
-------
That's a parabola. Since the coefficient of the x^2 term is +, it opens upward and has a minumum. The minimum is the vertex.
The vertex is on the LOS (Line of Symmetry).
The equation for the line of symmetry is x = -b/2a
x = 64/10 = 6.4
-------
2x + y = 16
--> y = 3.2
-----------
The answer is 6.4 & 3.2